OpenStudy (anonymous):

Find the inverse Laplace transform of F(s)=(2s+2)/(s^2+2s+5).

OpenStudy (anonymous):

Did you even try?

OpenStudy (anonymous):

I did, after completing the square, I got (2s+2)/(s+1)^2+4). But I don't know what to do after that.

OpenStudy (anonymous):

Well I do not know this tbh, I can send you a link to were some one has gotten the answer for this if you want it?

OpenStudy (anonymous):

Yes, please.

OpenStudy (anonymous):

I hope this is what you are looking for!

OpenStudy (anonymous):

It was somewhat helpful.

OpenStudy (loser66):

I think I can help you now \[F(s) = \frac{2(s-(-1))}{(s-(-1))^2+2^2}\] \[L^{-1}F(s) = 2 e^{-t}cos2t\]

OpenStudy (anonymous):

But how did you get the final answer?

OpenStudy (loser66):

that the final answer, use Laplace transform table.

OpenStudy (anonymous):

I got 2s/(s^2+2^2), but is there another way to find the final answer?

OpenStudy (loser66):

it's not that, friend. Your denominator doesn't lead to your answer. It's s^2 +2s +5, how can it be = x^2 +4?

OpenStudy (loser66):

*s not x

OpenStudy (anonymous):

No, I got s=s+1 from 2(s+1)/(s+1)^2+4, 2s/(s^2+4)=2s/(s^2+2^2)

OpenStudy (loser66):

I think you misunderstand the concept. you have f(x) you have to find Laplace inverse of that f(x) . You try to make it looked like some of the formula in Laplace transform table to get f(x)

OpenStudy (anonymous):

So what's the formula for this problem?

OpenStudy (loser66):

wait, I search and send you

OpenStudy (loser66):

look at the line 8, it talks about L^(-1) , the right column is what you have, it will lead to the left column answer. That's what I got from your problem. Our responsibility is try to convert the problem into one of the form of the right column, then use this table to find out the answer. Your prof didn't tell you that?

OpenStudy (anonymous):

Thanks for the site. Unfortunately, my professor sucks, that's why I need so much help in here.

OpenStudy (loser66):

ok.